# arXiv:2603.03170v1[math.AP] 03 Mar 2026

### v1

@@ 1. Introduction
The aim of this paper is to drop the traditional regularity assumptions in (1.1) and prove that the corresponding Cauchy problem is well-posed in the very weak sense. The notion of very weak solution has been introduced in the context of hyperbolic equations in [12] to deal with discontinuous coefficients and employed in related work on hyperbolic equations and systems with multiplicities [8, 9]. The same ideas have been recently applied in the context of Schrödinger type equations with singular coefficients in [2, 3]. In a nutshell, since the presence of distributional coefficients might not allow a meaningful definition of the operator $P$, our approach is to replace $P$ with a family of regularised operators $(P_{\varepsilon})_{\varepsilon}$ where $\varepsilon\to 0$. This is done via convolution with mollifiers of the type $\varphi_{\omega(\varepsilon)}(x)=\varepsilon^{-n}\varphi(x/\omega(\varepsilon))$ where $\omega$ is a regularising net tending to 0. Our aim is to identify a suitable regularisation method (mollifier and scale) such that the regularised Cauchy problem
@@
where $a_{ij}(x)=c_{ij}+\tilde{a}_{ij}(x)$, $c_{ij}\in\mathbb{R}$ and $\tilde{a}_{ij}\in W^{1,\infty}(\mathbb{R}^{n})$ is real valued, for all $i,j=1,\dots,n$. We identify a set of hypotheses on the regularised operator $A_{\varepsilon}$ that allow to generalise the Doi’s lemma for ultrahyperbolic operators with regular coefficients to this singular context. This is fundamental step in the proof of our well-posedness result. The full statement of the problem (involving lower order terms as well) and the proof of existence and uniquenss of a $H^{\infty}$-very weak solution are the contents of Section 4 and 5. Our proof method makes use of smoothing estimates as in [6, 7] combined with careful approximation techniques that states clearly the dependence on the regularising scale. We conclude the paper with Section 6 where consistency with the classical result, i.e., when the equation coefficients are smooth then every weak solution will converge to the classical solution as $\varepsilon\to 0$.
@@ 2.1. Basics of pseudodifferential calculus and Sobolev spaces
In order to give our well-posedness result for the problem (3.7) we will work on Sobolev spaces, defined through $\Lambda^{s}=\mathrm{Op}(\langle\xi\rangle^{s})$, as
@@
(2.6) $H^{\infty}(\mathbb{R}^{n}):=\bigcap_{s\in\mathbb{R}}H^{s}(\mathbb{R}^{n}),\quad H^{-\infty}(\mathbb{R}^{n}):=\bigcup_{s\in\mathbb{R}}H^{s}(\mathbb{R}^{n}).$
We conclude this subsection by summarizing the following fundamental results of the standard pseudodifferential calculus (cf. Section 2.1 in [11]) that we will use in the next sections.
@@
Let $a\in S^{m}$, with $m\in\mathbb{R}$, and let $s\in\mathbb{R}$. Then $\mathrm{Op}(a)$ extends to a bounded linear operator from $H^{s+m}(\mathbb{R}^{n})$ to $H^{s}(\mathbb{R}^{n})$ and there exist $k=k(n,m,s)\in\mathbb{N}$ and $C=C(n,m,s)>0$ such that
@@
where $r_{m_{1}+m_{2}-1}\in S^{m_{1}+m_{2}-1}$ and, for each $k\in\mathbb{N}$, there exists $k_{1}\in\mathbb{N}$ and $c_{1}>0$, depending on $m_{1},m_{2},k$, such that
@@
with $r_{m_{1}+m_{2}-2}\in S^{m_{1}+m_{2}-2}$ and, for each $k\in\mathbb{N}$, there exists $k_{2}\in\mathbb{N}$ and $c_{2}>0$, depending once again on $m_{1},m_{2},k$, such that
@@
with $r_{m-1}\in S^{m-1},$ and for each $k\in\mathbb{N}$ there exists $k^{\prime}\in\mathbb{N}$ and $c^{\prime}>0$, depending on $m$ and $k$, such that
@@
Then, there exist $k=k(n)\in\mathbb{N}$ and $C=C(n,R)>0$ such that
@@ 2.2. Regularisations via convolution with a mollifier
As before, we denote by $C>0$ a constant, possibly changing from line to line, independent of $\varepsilon$. Since $\Lambda^{s+\ell}=\mathscr{F}_{\xi\rightarrow x}^{-1}\langle\xi\rangle^{s+\ell}\mathscr{F}_{y\rightarrow\xi}$, by Plancherel’s theorem, we have
@@ 2.3. Sobolev moderate and negligible nets
We recall here the notion of $H^{s}$ and $H^{\infty}$ moderate nets of functions (cf. [12], [3],[2]) that, in the next sections, will lead to the notion of $H^{s}$ and $H^{\infty}$ very weak solution of the problem (1.1). Furthermore, we also recall the notion of negligible net, that will be useful to study the uniqueness result for very weak solutions. In what follows, we consider $T>0$ and $s\in\mathbb{R}$ as fixed, and we give the next definitions.
@@
Finally, if is useful to recall the definition of negligible nets.
@@ 3.1. Regularisation of singular ultrahyperbolic operators
Recall that the space $W^{1,\infty}(\mathbb{R}^{n})$ essentially coincides with the space of bounded Lipschitz functions. Therefore, one of the interests of our theory is that it allows us to establish well-posedness results for Cauchy problems induced by variable-coefficient operators whose principal part is obtained as a bounded Lipschitz perturbation of a constant-coefficient operator and, as we will see below, with much more singular lower-order terms.
The purpose now is to construct a net of operator $(A_{\varepsilon})_{\varepsilon}$ associated with $A$ obtained by a regularisation argument via convolution by a mollifier $\varphi_{\omega(\varepsilon)}$ defined (as in (2.7)) by
@@
Specifically, for all $i,j=1,\dots,n$, we define the the $\varepsilon$-regularised coefficients of $A$, defined in (3.10), by
@@
We thus define, for $\varepsilon\in(0,1]$, the $\varepsilon$-regularised operator by
@@
and denote by $\mathcal{A}_{\varepsilon}(x)=(a_{ij,\varepsilon}(x))_{i,j=1}^{n}$ the matrix of $\varepsilon$-regularised coefficients. We denote the set of regularised operators as $(A_{\varepsilon})_{\varepsilon}$ and we call it the net of regularised operators associated with $A$.
@@
Let $A_{\varepsilon}$ be the operator defined in (3.13), with $\varepsilon\in(0,1]$. Then, $A_{\varepsilon}=\mathrm{Op}(a_{\varepsilon})$ with the symbol $a_{\varepsilon}(x,\xi)$ fulfilling the following statement: for all $\alpha,\beta\in\mathbb{N}_{0}^{n}$ there exists a constant $C_{\alpha,\beta}>0$ such that
@@
with $a_{ij,\varepsilon}$ defined as in (3.12). Then, since $\tilde{a}_{ij}\in W^{1,\infty}(\mathbb{R}^{n})$ for all $i,j=1,\dots,n$, by (2.9) we obtain that for all $\alpha,\beta\in\mathbb{N}_{0}^{n}$ there exists $C_{\alpha,\beta}^{\prime}>0$ such that
@@
As anticipated above, let us emphasise that, once these conditions are required on the net $(A_{\varepsilon})_{\varepsilon}$, the operator $A$ can be interpreted as the "singular-coefficients" version of the ultrahyperbolic operators studied in [7] and [11]. We conclude this subsection with a few comments and motivations for the conditions we assumed.
@@ 3.2. Doi’s Lemma for ultrahyperbolic operators with singular coefficients
Let $A$ be an operator as in (3.10), with coefficients of the form (3.11) and denote by $(A_{\varepsilon})_{\varepsilon}$ the net of regularised operators. If (H1), (H2) and (H3) hold, then there exists a net of real-valued functions $(q_{\varepsilon})_{\varepsilon}$, with $q_{\varepsilon}\in C^{\infty}(\mathbb{R}^{n}\times\mathbb{R}^{n})$ that satisfies:
@@
for all $(x,\xi)\in\mathbb{R}^{n}\times\mathbb{R}^{n}$, where recall the Hamilton vector field $H_{a_{2},\varepsilon}$ is defined in (2.4).
@@
Let $A$ be an operator as in (3.10) with coefficients of the form (3.11), and denote by $(A_{\varepsilon})_{\varepsilon}$ the net of regularised operators. Moreover, assume that there exists a net of real-valued functions $(q_{\varepsilon})_{\varepsilon}$, with $q_{\varepsilon}\in C^{\infty}(\mathbb{R}^{n}\times\mathbb{R}^{n})$ that satisfies:
@@
Then, there exists a net of symbols $(d_{\varepsilon})_{\varepsilon}$ such that for all $\alpha,\beta\in\mathbb{N}_{0}^{n}$ and for each $\varepsilon\in(0,1]$
(3.27) $|\partial_{x}^{\beta}\partial_{\xi}^{\alpha}d_{\varepsilon}(x,\xi)|\leq C_{\alpha,\beta}\omega(\varepsilon)^{-|\beta|+1}\langle\xi\rangle^{-|\alpha|},\quad\forall(x,\xi)\in\mathbb{R}^{n}\times\mathbb{R}^{n},$
for some $C_{\alpha,\beta}>0$ and for all $\varepsilon\in(0,1]$
(3.28) $H_{a_{2,\varepsilon}}d_{\varepsilon}(x,\xi)\geq\langle x\rangle^{-N}|\xi|-C,\quad\mathrm{for}\ \mathrm{all}\ (x,\xi)\in\mathbb{R}^{n}\times\mathbb{R}^{n},$
@@ 4.1. Statement of the problem
and we work under the following additional hypotheses on the "first-order" and "zero-order" parts of the operator $P$.
@@
Note that $B_{\varepsilon}=\mathrm{Op}(b_{\varepsilon})$, where for each $\varepsilon\in(0,1]$, with the same notation of above, one has
@@
Henceforth, we sometimes say that $(P_{\varepsilon})_{\varepsilon}$ satisfy condition (H1), (H2), (H3), (H4), (H5), which means that the nets of coefficients of $(P_{\varepsilon})_{\varepsilon}$ we consider satisfy such conditions.
@@ 4.2. Regularised Problem
The proof of Theorem 4.4 follows the same pattern as Lemma 4.1 in [7] (cf. [11], [6]). We refer the reader to those works for further details.
@@
We now use this theorem to prove the existence of an $H^{\infty}$-weak solution of the problem (1.1).
@@ 5.1. Existence
$\begin{split}\sup_{0\leq t\leq T}\|u_{\varepsilon}(t,\cdot)\|_{s}&\leq C_{2}\omega(\varepsilon)^{-k_{2}}e^{TC_{1}\omega(\varepsilon)^{-k_{1}}}\left(\|u_{0,\varepsilon}\|_{s}+\int_{0}^{T}\|f_{\varepsilon}(t,\cdot)\|_{s}dt\right)\\ &\leq C_{2}^{\prime}e^{TC_{1}^{\prime}\omega(\varepsilon)^{-k_{1}}}\varepsilon^{-N_{f}-N_{u_{0}}}.\end{split}$
@@
we have the $H^{\infty}$-moderateness of $(u_{\varepsilon})_{\varepsilon}$, and then the existence of an $H^{\infty}$-weak solution, follows. ∎
@@ 6. Consistency with the classical theory
Hence, we have that $u-u_{\varepsilon}$ solves,
@@
and this proves the consistency of our approach with the classical theory.

### v2

# Schrödinger ultrahyperbolic equations with singular coefficients

@@ Abstract.
Key words and phrases:
Schrödinger operator, Ultrahyperbolic operator, singular coefficients, regularisation
2020 Mathematics Subject Classification:
Primary 35J10; Secondary 35B65, 35D99;
The authors were supported by the EPSRC grant EP/V005529/2. The second author is member of the Research Group GNAMPA of INdAM
On behalf of all authors, the corresponding author states that there is no conflict of interest. This manuscript has no associated data.
@@ 1. Introduction
The aim of this paper is to drop the traditional regularity assumptions in (1.1) and prove that the corresponding Cauchy problem is well-posed in the very weak sense. The notion of very weak solution has been introduced in the context of hyperbolic equations in [13] to deal with discontinuous coefficients and employed in related work on hyperbolic equations and systems with multiplicities [6, 7, 12, 14, 17]. The same ideas have been recently applied in the context of Schrödinger type equations with singular coefficients in [2, 3, 4, 18, 19]. In a nutshell, since the presence of distributional coefficients might not allow a meaningful definition of the operator $P$, our approach is to replace $P$ with a family of regularised operators $(P_{\varepsilon})_{\varepsilon}$ where $\varepsilon\to 0$. This is done via convolution with mollifiers of the type $\varphi_{\omega(\varepsilon)}(x)=\varepsilon^{-n}\varphi(x/\omega(\varepsilon))$ where $\omega$ is a regularising net tending to 0. Our aim is to identify a suitable regularisation method (mollifier and scale) such that the regularised Cauchy problem
@@
where $a_{ij}(x)=c_{ij}+\tilde{a}_{ij}(x)$, $c_{ij}\in\mathbb{R}$ and $\tilde{a}_{ij}\in W^{1,\infty}(\mathbb{R}^{n})$ is real valued, for all $i,j=1,\dots,n$. We identify a set of hypotheses on the regularised operator $A_{\varepsilon}$ that allow to generalise the Doi’s lemma for ultrahyperbolic operators with regular coefficients to this singular context. This is fundamental step in the proof of our well-posedness result. The full statement of the problem (involving lower order terms as well) and the proof of existence and uniquenss of a $H^{\infty}$-very weak solution are the contents of Section 4 and 5. Our proof method makes use of smoothing estimates as in [9, 10, 11] combined with careful approximation techniques that states clearly the dependence on the regularising scale. We conclude the paper with Section 6 where consistency with the classical result, i.e., when the equation coefficients are smooth then every weak solution will converge to the classical solution as $\varepsilon\to 0$.
@@ 2.1. Basics of pseudodifferential calculus and Sobolev spaces
Our well-posedness result will be formulated in the context of Sobolev spaces, defined through $\Lambda^{s}=\mathrm{Op}(\langle\xi\rangle^{s})$, as
@@
(2.6) $H^{\infty}(\mathbb{R}^{n}):=\bigcap_{s\in\mathbb{R}}H^{s}(\mathbb{R}^{n}),\quad H^{-\infty}(\mathbb{R}^{n}):=\bigcup_{s\in\mathbb{R}}H^{s}(\mathbb{R}^{n}),$
respectively. We conclude this subsection by summarizing the following fundamental results of the standard pseudodifferential calculus (cf. Section 2.1 in [16]) that we will use in the next sections.
@@
Let $a\in S^{m}$, with $m\in\mathbb{R}$, and let $s\in\mathbb{R}$. Then $\mathrm{Op}(a)$ extends to a bounded linear operator from $H^{s+m}(\mathbb{R}^{n})$ to $H^{s}(\mathbb{R}^{n})$ and there exist $k=k(n,m,s)\in\mathbb{N}_{0}$ and $C=C(n,m,s)>0$ such that
@@
where $r_{m_{1}+m_{2}-1}\in S^{m_{1}+m_{2}-1}$ and, for each $k\in\mathbb{N}_{0}$, there exists $k_{1}\in\mathbb{N}_{0}$ and $c_{1}>0$, depending on $m_{1},m_{2},k$, such that
@@
with $r_{m_{1}+m_{2}-2}\in S^{m_{1}+m_{2}-2}$ and, for each $k\in\mathbb{N}_{0}$, there exists $k_{2}\in\mathbb{N}_{0}$ and $c_{2}>0$, depending once again on $m_{1},m_{2},k$, such that
@@
with $r_{m-1}\in S^{m-1},$ and for each $k\in\mathbb{N}_{0}$ there exists $k^{\prime}\in\mathbb{N}_{0}$ and $c^{\prime}>0$, depending on $m$ and $k$, such that
@@
Then, there exist $k=k(n)\in\mathbb{N}_{0}$ and $C=C(n,R)>0$ such that
@@ 2.2. Regularisations via convolution with a mollifier
As before, we denote by $C$ a positive constant, possibly changing from line to line, independent of $\varepsilon$. Since $\Lambda^{s+\ell}=\mathscr{F}_{\xi\rightarrow x}^{-1}\langle\xi\rangle^{s+\ell}\mathscr{F}_{y\rightarrow\xi}$, by Plancherel’s theorem, we have
@@ 2.3. Sobolev moderate and negligible nets
We recall here the notions of $H^{s}$- and $H^{\infty}$-moderate nets of functions (cf. [13], [4],[3]) that, in the next sections, will lead to the notion of $H^{s}$- and $H^{\infty}$-very weak solution of the problem (1.1). Furthermore, we also recall the notion of negligible net, that will be useful to treat uniqueness in the context of very weak solutions. In what follows, we consider $T>0$ and $s\in\mathbb{R}$ as fixed.
@@
We conclude this section with the definition of negligible nets.
@@ 3.1. Regularisation of singular ultrahyperbolic operators
We recall that the space $W^{1,\infty}(\mathbb{R}^{n})$ essentially coincides with the space of bounded Lipschitz functions. Therefore, one of the interests of our theory is that it allows us to establish well-posedness results for Cauchy problems induced by variable-coefficient operators whose principal part is obtained as a bounded Lipschitz perturbation of a constant-coefficient operator and, as we will see below, with much more singular lower-order terms.
Our purpose is now to construct a net of operator $(A_{\varepsilon})_{\varepsilon}$ associated with $A$ obtained via a regularisation argument. In detail, we will comvolve with a mollifier $\varphi_{\omega(\varepsilon)}$ defined (as in (2.7)) by
@@
For all $i,j=1,\dots,n$, we define the the $\varepsilon$-regularised coefficients of $A$, defined in (3.10), as
@@
We thus define, for $\varepsilon\in(0,1]$, the $\varepsilon$-regularised operator
@@
and denote by $\mathcal{A}_{\varepsilon}(x)=(a_{ij,\varepsilon}(x))_{i,j=1}^{n}$ the matrix of $\varepsilon$-regularised coefficients. Finally, $(A_{\varepsilon})_{\varepsilon}$ is the net of regularised operators associated with $A$.
@@
Let $A_{\varepsilon}$ be the operator defined in (3.13), with $\varepsilon\in(0,1]$. Then, $A_{\varepsilon}=\mathrm{Op}(a_{\varepsilon})$ is an operator with symbol $a_{\varepsilon}(x,\xi)$ fulfilling the following statement: for all $\alpha,\beta\in\mathbb{N}_{0}^{n}$ there exists a constant $C_{\alpha,\beta}>0$ such that
@@
and $a_{ij,\varepsilon}$ defined as in (3.12). Then, since $\tilde{a}_{ij}\in W^{1,\infty}(\mathbb{R}^{n})$ for all $i,j=1,\dots,n$, by (2.9) we obtain that for all $\alpha,\beta\in\mathbb{N}_{0}^{n}$ there exists $C_{\alpha,\beta}^{\prime}>0$ such that
@@
As anticipated above, let us emphasise that, once these conditions are required on the net $(A_{\varepsilon})_{\varepsilon}$, the operator $A$ can be interpreted as the "singular-coefficients" version of the ultrahyperbolic operators studied in [11] and [16]. We conclude this subsection with a few comments and motivations related to the hypotheses (H1)-(H3).
@@ 3.2. Doi’s Lemma for ultrahyperbolic operators with singular coefficients
Let $A$ be an operator as in (3.10), with coefficients of the form (3.11) and let $(A_{\varepsilon})_{\varepsilon}$ be the corresponding net of regularised operators. If (H1), (H2) and (H3) hold, then there exists a net of real-valued functions $(q_{\varepsilon})_{\varepsilon}$, with $q_{\varepsilon}\in C^{\infty}(\mathbb{R}^{n}\times\mathbb{R}^{n})$ that satisfies the following:
@@
for all $(x,\xi)\in\mathbb{R}^{n}\times\mathbb{R}^{n}$, where $H_{a_{2},\varepsilon}$ is the Hamilton vector field defined in (2.4).
@@
Let $A$ be an operator as in (3.10) with coefficients of the form (3.11), and let $(A_{\varepsilon})_{\varepsilon}$ be the corresponding net of regularised operators. Moreover, assume that there exists a net of real-valued functions $(q_{\varepsilon})_{\varepsilon}$, with $q_{\varepsilon}\in C^{\infty}(\mathbb{R}^{n}\times\mathbb{R}^{n})$ that satisfies the following statements:
@@
Then, there exists a net of symbols $(d_{\varepsilon})_{\varepsilon}$ such that for all $\alpha,\beta\in\mathbb{N}_{0}^{n}$
(3.27) $|\partial_{x}^{\beta}\partial_{\xi}^{\alpha}d_{\varepsilon}(x,\xi)|\leq C_{\alpha,\beta}\omega(\varepsilon)^{-|\beta|+1}\langle\xi\rangle^{-|\alpha|},\quad\forall(x,\xi)\in\mathbb{R}^{n}\times\mathbb{R}^{n},\,\forall\varepsilon\in(0,1]$
for some $C_{\alpha,\beta}>0$ and
(3.28) $H_{a_{2,\varepsilon}}d_{\varepsilon}(x,\xi)\geq\langle x\rangle^{-N}|\xi|-C,\quad\forall(x,\xi)\in\mathbb{R}^{n}\times\mathbb{R}^{n},\,\forall\varepsilon\in(0,1],$
@@ 4.1. Statement of the problem
and we work under the following additional hypotheses on the "first-order" and "zero-order" parts of the operator $P$. We assume:
@@
Note that $B_{\varepsilon}=\mathrm{Op}(b_{\varepsilon})$, where for each $\varepsilon\in(0,1]$, with the same notation as above, one has
@@
Henceforth, we say that $(P_{\varepsilon})_{\varepsilon}$ satisfies condition (H1), (H2), (H3), (H4), (H5), which means that the nets of coefficients of $(P_{\varepsilon})_{\varepsilon}$ we consider satisfy such conditions.
@@ 4.2. Regularised Problem
The proof of Theorem 4.4 follows the same pattern as Lemma 4.1 in [11] (cf. [16], [10]). We refer the reader to [11] for further details.
@@
We are now ready to employ this theorem to prove the existence of an $H^{\infty}$-weak solution of the Cauchy problem (1.1).
@@ 5.1. Existence
$\begin{split}\sup_{0\leq t\leq T}\|u_{\varepsilon}(t,\cdot)\|_{s}&\leq C_{2}e^{TC_{1}\omega(\varepsilon)^{-k_{1}}}\left(\|u_{0,\varepsilon}\|_{s}+\int_{0}^{T}\|f_{\varepsilon}(t,\cdot)\|_{s}dt\right)\\ &\leq C_{2}^{\prime}e^{TC_{1}\omega(\varepsilon)^{-k_{1}}}\varepsilon^{-N_{f}-N_{u_{0}}}.\end{split}$
@@
we have the $H^{\infty}$-moderateness of $(u_{\varepsilon})_{\varepsilon}$, and then the existence of an $H^{\infty}$-weak solution, follows. Note that the scale $\omega$ has been chosen to be independent of $k_{1}$ and therefore independent of the Sobolev order $s$, leading to $H^{\infty}$-moderateness. ∎
@@ 6. Consistency with the classical theory
Proposition 6.6.
Under the hypotheses above any very weak solution $u_{\varepsilon}$ converges to the classical solution $u$ in $C([0,T];H^{\infty}(\mathbb{R}^{n}))$.
Proof.
Note that $u-u_{\varepsilon}$ solves,
@@
and this proves the consistency of our approach with the classical theory. ∎